# Gaming on Coincident Peak Shaving: Equilibrium and Strategic Behavior

### v4

@@ III-B Nash Equilibrium
In this section, we study the NEs of the two-agent two-period CP games in all conditions as described in Proposition 4. The main theorem is as follows
@@
NEs in two-agent two-period CP game. Given critical points $r_{\mathrm{x}},r_{\mathrm{y}}$ and balance points $b_{\mathrm{x}},b_{\mathrm{y}},b$ as defined in Definition 2, the unique pure-strategy NE $(x^{*},y^{*})$ as defined in Definition 1 of the two-agent two-period CP game $G$ satisfy one of the following:
@@
    Concave CP game.
    $x^{*}$ $=-r_{\mathrm{x}},y^{*}=-r_{\mathrm{y}},$ $\text{if }b_{\mathrm{x}}<-r_{\mathrm{x}},b_{\mathrm{y}}<-r_{\mathrm{y}}-r_{\mathrm{x}}-b_{\mathrm{x}};$ (7a)
    $x^{*}$ $=r_{\mathrm{x}},y^{*}=r_{\mathrm{y}},$ $\text{if }b_{\mathrm{x}}>r_{\mathrm{x}},b_{\mathrm{y}}>r_{\mathrm{y}}+r_{\mathrm{x}}-b_{\mathrm{x}};$ (7b)
@@
    Quasiconcave CP game.
    $x^{*}$ $=b_{\mathrm{x}},y^{*}=b_{\mathrm{y}},$ $\text{if }-r_{\mathrm{x}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}},-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq r_{\mathrm{y}};$ (7c)
@@
    Non-concave CP game.
    $x^{*}$ $=-r_{\mathrm{x}},y^{*}=b+r_{\mathrm{x}},$ if $\text{if }\{-r_{\mathrm{y}}-r_{\mathrm{x}}\leq b\leq r_{\mathrm{y}}-r_{\mathrm{x}}\}\cap\{b_{\mathrm{x}}<-r_{\mathrm{x}}\cup b_{\mathrm{y}}>r_{\mathrm{y}}\}$ (7d)
    $x^{*}$ $=r_{\mathrm{x}},y^{*}=b-r_{\mathrm{x}},$ if $\text{if }\{-r_{\mathrm{y}}+r_{\mathrm{x}}\leq b\leq r_{\mathrm{y}}+r_{\mathrm{x}}\}\cap\{b_{\mathrm{x}}>r_{\mathrm{x}}\cup b_{\mathrm{y}}<-r_{\mathrm{y}}\}$ (7e)
@@
We comprehensively analyze the NE in all possible conditions by separating two cases according to the system CP period determined by $S_{\mathrm{b},1},S_{\mathrm{b},2}$, and further separate the cases according to agents’ own baseline demand $X_{1},X_{2},Y_{1},Y_{2}$, then study whether both agents are capable, one of them is capable, and none of them is capable. The reason is that agents’ best response is to shift demand away from the CP period when their demand is higher in the CP period, to avoid CP charges, or shift demand toward the CP period when their demand is lower in the CP period, to maintain the CP period. However, the shifting capability is determined by their parameters, i.e., the relationship between the critical point $r_{\mathrm{x}}$ and the agent balance point $b_{\mathrm{x}}$. We analyze each condition and show corresponding NE solutions. The details are provided in the appendix. ∎
This theorem shows the NEs under each concave, quasiconcave, and non-concave CP game type. Note that the condition described in this theorem aligns with the conditions in Proposition 4, and we write out individual agent’s conditions along with the NE solutions. Basically, the concave CP game includes two cases (7a) and (7b), both are aligned with (6a) and (6b) in Proposition 4. The condition of quasiconcave CP game (7c) is the same to (6c) in Proposition 4. The non-concave CP game conditions (7d) and (7e) are also aligned with (6d) in Proposition 4.
@@
The NE also describes the relationship of gaming agents as fully cooperative (7a), (7b), fully competitive (7c), and mixed competitive and cooperative (7d), (7e). Specifically, When they are in a fully cooperative relationship, they cooperate to shift the demand away from the CP period. Once they can change the CP period together, they enter a competitive relationship where they compete with each other to change the CP period, but one agent’s decision is limited by their shifting capability. When all agents are capable, they are in a fully competitive relationship that can balance their demand regardless of opponents’ strategy. Indeed, if both agents want to be in a mixed relationship, they must be ’asymmetrical,’ where their shifting penalty parameters should be different so that they have different shifting limitations. Otherwise, they are either fully cooperative or fully competitive.
@@
We also show in practice, under the quasiconcave and concave game conditions, agents share the market equally, i.e., their strategy only depends on their parameters, regardless of the opponent’s strategy. While in the non-concave game, the agents’ strategy depends on the opponent’s strategy, and the agent with $y=b\mp r_{\mathrm{x}}$ is more flexible and dominates the other inflexible agent with $x=\pm r_{\mathrm{x}}$ and benefit more through the competition. We denote agents’ flexibility as the demand shifting capability, either $x^{*}=\min\{r_{\mathrm{x}},b_{\mathrm{x}}\}$ or $x^{*}=\max\{-r_{\mathrm{x}},b_{\mathrm{x}}\}$. Here, $x^{*}$ depends on the game type and is influenced by the shifting penalty parameter $\alpha_{\mathrm{x}}$ and $b_{\mathrm{x}}$. Thus, flexibility is determined by both $\alpha$ and the baseline demand conditions $X_{1},X_{2},Y_{1},Y_{2}$.
@@ IV Equilibrium stability and algorithm convergence
The CP game system is a switched dynamics system due to the indicator function in the model (1). In this section, we first analyze the system stability and then develop a solution algorithm to reach the stable (equilibrium) point.
@@ IV-A Equilibrium stability
There are two system dynamics corresponding to periods 1 and 2, separated by the indicator function and a switching logic between these two periods. According to [Rosen, 1965], we consider a reasonable dynamic model in each period in which each agent changes his strategy following the gradient direction with respect to his strategy of his payoff function, then each agent’s payoff will increase given all other agents’ strategies. Denote the dynamic time index with $k$, and the gradient as $F_{1}(x,y)=[\nabla_{\mathrm{x}}f_{\mathrm{x}}(x,y),\nabla_{\mathrm{y}}f_{\mathrm{y}}(x,y)]^{T}$ with total differential operator $\nabla$, the dynamics of period 1 ($x,y\in\mathcal{X}_{\mathrm{cp}1}$) is
@@
We then denote the CP game system with the switching logic as
@@
Our goal is to prove the overall CP game system with the switching logic is global uniform asymptotically stable. According to the NE described in Theorem 5, we call $(x^{*},y^{*})$ as the equilibrium point of the CP game system (9) if one of the following is satisfied,
@@
$F_{1}(x^{*},y^{*})$ $=0,S_{\mathrm{b},1}\geq S_{\mathrm{b},2},$ (10b)
$F_{2}(x^{*},y^{*})$ $=0,S_{\mathrm{b},1}<S_{\mathrm{b},2}.$ (10c)
Among them, the first condition (10a) corresponds to non-concave CP game (7d) and (7e), and quasiconcave CP game (7c), where both periods have the same CP charges and shifting penalty because the system demand is balanced in the two periods at NE and the shifting penalty is symmetric in the two periods. The second and third (10b) and (10c) conditions correspond to the concave CP game (7a) and (7b), where the NE is obtained when the gradient $F_{1}$ or $F_{2}$ reaches zero. Then, we introduce the main stability results.
@@
Global stability of equilibrium point. The CP game system (9) is global uniform asymptotically stable in $\mathcal{X}_{\mathrm{s}}$, where
@@
i.e., for every starting point $(x,y)\in\mathcal{X}_{\mathrm{s}}$, the solution $(x(k),y(k))$ to the CP game system (9) converges to an equilibrium point $(x^{*},y^{*})\in\mathcal{X}_{\mathrm{s}}$ as $k\rightarrow\infty$, where $(x^{*},y^{*})$ is the equilibrium point satisfy (10).
@@
The switched dynamics system property makes the stability analysis non-trivial. We first show that the dynamical systems within each period described by (8a) and (8b) are asymptotically stable in $\mathcal{X}$. Then, we include the switching logic and derive multiple Lyapunov functions with continuous properties in the switching surface to show the CP game system (9) is global uniform asymptotically stable in $\mathcal{X}_{\mathrm{s}}$. After that, we show the stable point is obtained with $F_{1}=0$ or $F_{2}=0$ under the condition of concave CP game, and with $f_{\mathrm{x},1}+f_{\mathrm{y},1}=f_{\mathrm{x},2}+f_{\mathrm{y},2}$ under the condition of quasiconcave or non-concave CP game. From Theorem 5, we also know the equilibrium is unique, which means the stable point is the equilibrium point. The detailed proof is provided in the appendix. ∎
This Theorem shows the global uniform asymptotically stability of our CP game system in the strategy set $\mathcal{X}_{\mathrm{s}}$. Combine with Theorem 5, the CP game $G$ can converge to the unique equilibrium points for the solution trajectory that within $\mathcal{X}_{\mathrm{s}}$. It is important to know that $\mathcal{X}_{\mathrm{s}}$ is reasonable in real operations. Specifically, the key component of $\mathcal{X}_{\mathrm{s}}$ is $\pi(S_{\mathrm{b},2}-x-y)$, which means when agents shift demand, the shifting amount remains within the system’s baseline limitation. This holds naturally and ensures practical system stability, as demand shifting cannot result in negative system demand, meaning agents cannot ”borrow” demand to shift beyond their available capacity.
@@ IV-B Algorithm to determine the equilibrium point
Given the finite difference approximation to the CP game system dynamics (9) with learning rate vector $\tau_{\mathrm{x},h},\tau_{\mathrm{y},h}$ for agent x and y, and gradient $F_{j}$
@@
where $\text{diag}(\cdot):\mathbb{R}^{2}\to\mathbb{R}^{2\cdot 2}$, $j$ is the switching signal taking the value 1 for $S_{1}(x,y)\geq S_{2}(x,y)$ and 2 for $S_{1}(x,y)<S_{2}(x,y)$. This then forms a gradient-based algorithm following the updating rule (12) to gradually reach the equilibrium point. Among them, the gradient is determined by the period on which the current solution lies, and the learning rate for agents x and y $\tau_{\mathrm{x},h},\tau_{\mathrm{y},h}$ are determined by its payoff function and gradient. Choosing a suitable learning rate is key to showing convergence performance, we thus provide the following Theorem.
@@
We use the backtracking line search method [Boyd and Vandenberghe, 2004] to calculate the learning rate, which is affected by the CP periods of the current and future steps. For the concave game, only one period could be the CP period during the entire solution process, according to the backtracking line search, the gradient $F_{1}$ or $F_{2}$ reduce gradually, and combined with Theorem 6, the gradient will reduce to zero and reach the equilibrium point satisfy (10b) or (10c).
@@
When switching happens, agents’ individual peak periods are different, and we analyze the objective function $-f_{\mathrm{x},1},-f_{\mathrm{x},2}$ change following the backtracking line search criteria by judging where the CP periods is in the current and next steps. Given a trajectory starts from period 1 at $h$, switch to period 2 at $h+1$, and back to period 1 at $h+2$, suppose agent x’s individual peak period is 1, we then show agent x updates its decision $x$ when switching from period 1 to 2, and agent y update its decision $y$ when switching from period 2 to 1, i.e., the $\tau_{\mathrm{y},h}=0,\tau_{\mathrm{x},h+1}=0$, correspondingly, agent x’s objective $-f_{\mathrm{x},1}$ decrease and agent y’s objective $-f_{\mathrm{y},2}$ decrease. The reason for the change is that their individual peak period aligns with the system’s CP period. Note that this learning rate is determined by their payoff functions and gradients, allowing the gradient for agent x to increase its decision variable while the gradient for agent y keeps the decision variable. By the same analysis, we know agent x’s objective $-f_{\mathrm{x},2}$ increase and agent y’s objective $-f_{\mathrm{y},1}$ increase in the trajectory starting from period 2, switched to period 1, and back to period 2. We then show these results still hold when the trajectory stays more steps in the system between two transitions. This proves the reduction of the distance between objective functions in the two periods. Combined with Theorem 6, the entire system finally will satisfy (10a) and reach the equilibrium point. The details are provided in the appendix. ∎
@@
This Theorem shows that the finite learning rate can be chosen in each step using the backtracking line search. Following this updating rule (12), combined with Theorem 6, a gradient-based algorithm can reach the equilibrium point. Specifically, the difference in each agent’s payoff function in the two periods reduces gradually, and depending on whether the game is concave or not, the difference can be reduced to zero or until one of the periods’ gradients reaches zero.
@@
As Theorem 6 shows, the system is asymptotically stable in $\mathcal{X}_{\mathrm{s}}$. Combined with Theorem 7, this confirms that the algorithm computes the NE as described in Theorem 5. When applying this game framework to a practical CP shaving system, the convergence process naturally reflects real agent interactions in the gaming process. Specifically, the system dynamics alternate between periods 1 and 2, while agents iteratively update their decisions based on the best response strategy. That is, given the opponents’ (aggregated) strategy, each agent determines an optimal response based on its own payoff. This iterative process follows the system dynamics and ultimately converges to the NE described in Theorem 5. Note that we don’t use a higher-order gradient descent method, such as Newton’s method because the fast updating with higher-order gradient information lets the solution in each period converge to each period’s stable point too fast to realize the converge on the overall switched system.

### v5

@@ III-B Nash Equilibrium
In this section, we study the NEs of the two-agent two-period CP shaving games in all conditions as described in Proposition 1. The main theorem is as follows.
@@
NEs in two-agent two-period CP shaving game. Given critical points $r_{\mathrm{x}},r_{\mathrm{y}}$ and balance points $b_{\mathrm{x}},b_{\mathrm{y}},b$ as defined in Definition 2, the unique pure-strategy NE $(x^{*},y^{*})$ as defined in Definition 1 of the two-agent two-period CP shaving game $G$ satisfy one of the following:
@@
    Concave CP shaving game.
    $x^{*}$ $=r_{\mathrm{x}},\ y^{*}=r_{\mathrm{y}},$ (7a)
@@
    Quasiconcave CP shaving game.
    $x^{*}$ $=b_{\mathrm{x}},\ y^{*}=b_{\mathrm{y}},$ (7b)
@@
    Non-concave CP shaving game.
    $x^{*}$ $=r_{\mathrm{x}},y^{*}=b-r_{\mathrm{x}},$ if $\text{if }\{0<b\leq r_{\mathrm{x}}+r_{\mathrm{y}}\}\cap\{b_{\mathrm{x}}>r_{\mathrm{x}}\},$
    $x^{*}$ $=b+r_{\mathrm{y}},y^{*}=-r_{\mathrm{y}},$ if $\text{if }\{0<b\leq r_{\mathrm{x}}+r_{\mathrm{y}}\}\cap\{b_{\mathrm{y}}<-r_{\mathrm{y}}\},$
    $x^{*}$ $=b-r_{\mathrm{y}},y^{*}=r_{\mathrm{y}},$ if $\text{if }\{0<b\leq r_{\mathrm{x}}+r_{\mathrm{y}}\}\cap\{b_{\mathrm{y}}>r_{\mathrm{y}}\}.$ (7c)
@@
We first provide a Lemma (Lemma 1) to establish the NE for the two-agent, two-period quasiconcave CP shaving game $G$, providing insight into the best response rationale. Then we separate two possible cases depends on agents’ baseline demand conditions, then study whether both agents are capable, one of them is capable, and none of them is capable. The reason is that agents’ best response is to shift demand away from the CP period when their individual demand is higher in the CP period, to avoid CP charges, or shift demand toward the CP period when their individual demand is lower in the CP period, to maintain the CP period. However, the shifting capability is determined by the relationship between the critical point $r_{\mathrm{x}}$ and the agent balance point $b_{\mathrm{x}}$. We analyze each condition and show corresponding NE solutions. The details are provided in the appendix. ∎
@@
This theorem shows the NEs and correspind conditions under each concave, quasiconcave, and non-concave CP shaving game type. The NE describes the relationship of gaming agents as fully cooperative (7a), fully competitive (7b), and mixed competitive and cooperative (7c). Specifically, When they are in a fully cooperative relationship, they cooperate to shift the demand away from the CP period. Once they can change the CP period together, they enter a competitive relationship where they compete with each other to change the CP period, but one agent’s decision is limited by their shifting capability. When all agents are capable, they are in a fully competitive relationship that can balance their demand regardless of opponents’ strategy. Indeed, if both agents want to be in a mixed relationship, they must be ’asymmetrical,’ where their shifting penalty parameters should be different so that they have different shifting limitations. Otherwise, they are either fully cooperative or fully competitive. Note that the condition described in this theorem aligns with the conditions in Proposition 1, specifically for non-concave game, we write out each condition conditions along with the NE solutions.
@@
We also demonstrate that, in practice, under quasiconcave and concave game conditions, agents share the market equally, meaning their strategies depend solely on their own parameters, independent of their opponent’s strategy. In contrast, in the non-concave game, an agent’s strategy is influenced by the opponent’s actions. For example, when one agent’s strategy follows $y=b\mp r_{\mathrm{x}}$ it tends to be more flexible and dominates the other, less flexible agent with $x=\pm r_{\mathrm{x}}$, gaining a competitive advantage and benefiting more. Notably, agents’ demand-shifting strategies $x^{*},y^{*}$ reflect their flexibility and are determined by the game type as well as the shifting penalty parameter $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}$ and baseline demand conditions $b_{\mathrm{x}},b_{\mathrm{y}}$ (depends on $X_{1},X_{2},Y_{1},Y_{2}$).
@@ IV Equilibrium stability and algorithm convergence
The CP shaving game system is a switched dynamics system due to the indicator function in the model (1). In this section, we first analyze the system stability and then develop a solution algorithm to reach the stable (equilibrium) point.
@@ IV-A Equilibrium stability
There are two system dynamics corresponding to periods 1 and 2, separated by the indicator function and a switching logic between these two periods. According to [Rosen,, 1965], we consider a reasonable dynamic model in each period in which each agent changes his strategy following the gradient direction with respect to his strategy of his payoff function, then each agent’s payoff will increase given all other agents’ strategies. Denote the dynamic time index with $k$, and the gradient as $F_{1}(x,y)=[\nabla_{\mathrm{x}}f_{\mathrm{x}}(x,y),\nabla_{\mathrm{y}}f_{\mathrm{y}}(x,y)]^{T}$ with total differential operator $\nabla$, the dynamics of period 1 ($x,y\in\mathcal{X}_{\mathrm{cp}1}$) is
@@
We then denote the CP shaving game system with the switching logic as
@@
Our goal is to prove the overall CP shaving game system with the switching logic is global uniform asymptotically stable [Liberzon,, 2003]. According to the NE described in Theorem 1, we call $(x^{*},y^{*})$ as the equilibrium point of the CP shaving game system (9) if one of the following is satisfied,
@@
$F_{2}(x^{*},y^{*})$ $=0.$ (10b)
Among them, the first condition (10a) corresponds to non-concave CP shaving game (7c) and quasiconcave CP shaving game (7b), where both periods have the same CP charges and shifting penalty because the system demand is balanced in the two periods at NE and the shifting penalty is symmetric in the two periods. The second (10b) conditions correspond to the concave CP shaving game (7a), where the NE is obtained when the gradient $F_{2}$ reaches zero. Then, we introduce the main stability results.
@@
Global stability of equilibrium point. The CP shaving game system (9) is global uniform asymptotically stable in $\mathcal{X}_{\mathrm{s}}$, where
@@
i.e., for every starting point $(x,y)\in\mathcal{X}_{\mathrm{s}}$, the solution $(x(k),y(k))$ to the CP shaving game system (9) converges to an equilibrium point $(x^{*},y^{*})\in\mathcal{X}_{\mathrm{s}}$ as $k\rightarrow\infty$, where $(x^{*},y^{*})$ is the equilibrium point satisfy (10).
@@
The switched dynamics system property makes the stability analysis non-trivial. We first show that the dynamical systems within each period described by (8a) and (8b) are asymptotically stable in $\mathcal{X}$. Then, we include the switching logic and derive multiple Lyapunov functions with continuous properties in the switching surface to show the CP shaving game system (9) is global uniform asymptotically stable in $\mathcal{X}_{\mathrm{s}}$ [Liberzon,, 2003]. After that, we show the stable point is obtained with $F_{2}=0$ under the condition of concave CP shaving game, and with $f_{\mathrm{x},1}+f_{\mathrm{y},1}=f_{\mathrm{x},2}+f_{\mathrm{y},2}$ under the condition of quasiconcave or non-concave CP shaving game. From Theorem 1, we also know the equilibrium is unique, which means the stable point is the equilibrium point. The detailed proof is provided in the appendix. ∎
This Theorem shows the global uniform asymptotically stability of our CP shaving game system in the strategy set $\mathcal{X}_{\mathrm{s}}$. Combine with Theorem 1, the CP shaving game $G$ can converge to the unique equilibrium points for the solution trajectory that within $\mathcal{X}_{\mathrm{s}}$. It is important to know that $\mathcal{X}_{\mathrm{s}}$ is reasonable in real operations. Specifically, the key component of $\mathcal{X}_{\mathrm{s}}$ is $\pi(S_{\mathrm{b},2}-x-y)$, which means when agents shift demand, the shifting amount remains within the system’s baseline limitation. This holds naturally and ensures practical system stability, as demand shifting cannot result in negative system demand, meaning agents cannot ”borrow” demand to shift beyond their available capacity.
@@ IV-B Algorithm to determine the equilibrium point
Given the finite difference approximation to the CP shaving game system dynamics (9) with learning rate vector $\tau_{\mathrm{x},h},\tau_{\mathrm{y},h}$ for agent x and y, as well as gradient $F_{j}$, we have
@@
where $\text{diag}(\cdot):\mathbb{R}^{2}\to\mathbb{R}^{2\cdot 2}$, $j$ is the switching signal taking the value 1 for $S_{1}(x,y)\geq S_{2}(x,y)$ and 2 for $S_{1}(x,y)<S_{2}(x,y)$. This then forms a gradient-based algorithm following the updating rule (12) to gradually reach the equilibrium point. Among them, the gradient is determined by the period on which the current solution lies, and the learning rate for agents x and y $\tau_{\mathrm{x},h},\tau_{\mathrm{y},h}$ are determined by its payoff function and gradient. Choosing a suitable learning rate is the key to showing convergence performance, we thus provide the following Theorem.
@@
We use the backtracking line search method [Boyd and Vandenberghe,, 2004] to calculate the learning rate, which is affected by the CP periods of the current and future steps. For the concave game, only the baseline CP period (2) could be the CP period during the entire solution process, according to the backtracking line search, the gradient $F_{2}$ reduce gradually, and combined with Theorem 2, the gradient will reduce to zero and reach the equilibrium point satisfy (10b).
@@
When switching happens, agents’ individual peak periods are different, and we analyze the objective function $-f_{\mathrm{x},1},-f_{\mathrm{x},2}$ change following the backtracking line search criteria by judging where the CP periods is in the current and next steps. Given a trajectory starts from period 1 at $h$, switch to period 2 at $h+1$, and back to period 1 at $h+2$, suppose agent x’s individual peak period is 1, we then show agent x updates its decision $x$ when switching from period 1 to 2, and agent y update its decision $y$ when switching from period 2 to 1, i.e., the $\tau_{\mathrm{y},h}=0,\tau_{\mathrm{x},h+1}=0$, correspondingly, agent x’s objective $-f_{\mathrm{x},1}$ decrease and agent y’s objective $-f_{\mathrm{y},2}$ decrease. The reason for the change is that their individual peak period aligns with the system’s CP period. Note that this learning rate is determined by their payoff functions and gradients, allowing the gradient for agent x to increase its decision variable while the gradient for agent y keeps the decision variable. By the same analysis, we know agent x’s objective $-f_{\mathrm{x},2}$ increase and agent y’s objective $-f_{\mathrm{y},1}$ increase in the trajectory starting from period 2, switched to period 1, and back to period 2. We then show these results still hold when the trajectory stays more steps in the period that between two transitions. This proves the reduction of the distance between the objective functions in the two periods. Combined with Theorem 2, the entire system finally will satisfy (10a) and reach the equilibrium point. The details are provided in the appendix. ∎
@@
This Theorem shows that the finite learning rate can be chosen in each step using the backtracking line search. Following this updating rule (12), combined with Theorem 2, a gradient-based algorithm can reach the equilibrium point. Specifically, the difference in each agent’s payoff function in the two periods reduces gradually, and depending on whether the game is concave or not, the difference can be reduced to zero or until baseline CP periods’ gradients reaches zero.
@@
As Theorem 2 shows, the system is asymptotically stable in $\mathcal{X}_{\mathrm{s}}$. Combined with Theorem 3, this confirms that the algorithm computes the NE as described in Theorem 1. When applying this game framework to a practical CP shaving system, the convergence process naturally reflects real agent interactions in the gaming process. Specifically, the system dynamics alternate between periods 1 and 2, while agents iteratively update their decisions based on the best response strategy. That is, given the opponents’ (aggregated) strategy, each agent determines an optimal response based on its own payoff. This iterative process follows the system dynamics and ultimately converges to the NE described in Theorem 1. Note that we don’t use a higher-order gradient descent method, such as Newton’s method, because the fast updating with higher-order gradient information lets the solution in each period converge to each period’s stable point too fast to realize the converge on the overall switched system.